OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. 1) square root of x+7 = x+1

OpenStudy (anonymous):

\[\sqrt{x+7}=x+1\]

OpenStudy (anonymous):

\[\sqrt{x + 7} = x + 1\] squaring both sides u get \[x+7 = (x + 1)^{2}\] this gives\[x + 7 = x ^{2} + 2x + 1\] now u just have to simplify and solve the resulting quadratic equation for x... if u need help, let me know

OpenStudy (anonymous):

I already tried doing that and i got x= -3, 2 but when i plug it in it doesn't work

OpenStudy (anonymous):

we get\[x ^{2} + x - 6 = 0\] x^2 + 3x - 2x - 6 = 0 x(x + 3) -2(x + 3) = 0 (x -2)(x + 3) = 0 so x = 2 or x = -3 and both are right I'll show thru drawing..

OpenStudy (anonymous):